# Learning-based Distributed Model Predictive Control using Multi-Agent Bayesian Optimization

### v1

@@ IV-B Multi-Agent Bayesian Optimization
In this paper, we aim to use a BO-based learning mechanism for adjusting the local MPC schemes in a distributed MPC framework. Therefore, we propose to leverage the BO method in the context of multi-agent systems introducing decomposable Multi-agent Bayesian Optimization (MABO).
@@
As observed in the parametric local MPC scheme (27), each local value function depends on the total parameters $\boldsymbol{\mathrm{\zeta}}:=\left\{\boldsymbol{\mathrm{\theta}}_{1},\ldots,\boldsymbol{\mathrm{\theta}}_{m}\right\}$ through the optimal multipliers $\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\left(\boldsymbol{\mathrm{\zeta}}\right)$ (each agent has full access to all multipliers using a communication network between local agents). Consequently, each local parametric policy $\boldsymbol{\mathrm{\pi}}_{i}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)$ delivered from (27) is also affected by the total parameters $\boldsymbol{\mathrm{\zeta}}$ and its neighbor policies $\boldsymbol{\mathrm{\pi}}_{j}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)$ so that the corresponding local closed-loop control performance reads as $J_{i}\left(\boldsymbol{\mathrm{\zeta}}\right):=J_{i}\left(\boldsymbol{\mathrm{\pi}}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)\right)$.
@@
$J_{i}^{N}\left(\boldsymbol{\mathrm{\zeta}}\right)=\frac{1}{N}\sum_{k=0}^{N}L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)\right)$ (33)
We then aim to implement a coordinated multi-agent learning problem by minimizing the global closed-loop performance as
@@
where $\rho>0$ is a constant penalty parameter and $\boldsymbol{\mathrm{\lambda}}=\left(\boldsymbol{\mathrm{\lambda}}_{1},\ldots,\boldsymbol{\mathrm{\lambda}}_{m}\right)$ denotes the set of dual variables (multipliers) associated to the equality constraint (35b). The local parameters are then obtained by the following local optimization problem
@@
where the coordinating variables $\boldsymbol{\mathrm{w}}^{k}:=(\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1},\boldsymbol{\mathrm{\lambda}}_{i}^{k})$ are updated by the following iterations
@@
To solve the problem (37), we then minimize the local acquisition function associated to $J_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i}\right)$ as
$\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}=\operatorname*{arg\,min}_{\boldsymbol{\mathrm{\zeta}}_{i}\in\mathcal{P}_{i}}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{w}}^{k}\right),$ (39)
where the penalty term $\Xi_{i}$ is computed as
$\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{w}}^{k}\right)=(\boldsymbol{\mathrm{\lambda}}_{i}^{k})^{\top}\boldsymbol{\mathrm{\zeta}}_{i}+\frac{\rho}{2}\left\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}\right\|^{2}.$ (40)
A multi-agent BO is described in Algorithm 1.
@@ 1  Input:
Coordinating variables $\boldsymbol{\mathrm{w}}^{k}:=(\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1},\boldsymbol{\mathrm{\lambda}}_{i}^{k})$
@@ 4  Output:
$\left(\boldsymbol{\mathrm{P}}_{i}^{k+1},\boldsymbol{\mathrm{Y}}_{i}^{k+1},\boldsymbol{\mathrm{w}}^{k+1}\right)=\textbf{{MABO}}\left(\boldsymbol{\mathrm{P}}_{i}^{k},\boldsymbol{\mathrm{Y}}_{i}^{k},\boldsymbol{\mathrm{w}}^{k}\right)$
@@
  3. Compute $\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}$ and $\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{w}}^{k}\right)$
  4. Find $\boldsymbol{\mathrm{\zeta}}_{i}^{\text{new}}=\operatorname*{arg\,min}_{\boldsymbol{\mathrm{\zeta}}_{i}\in\mathcal{P}_{i}}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{w}}^{k}\right)$
@@ 5. Evaluate $y_{i}^{\text{new}}=J_{i}^{N}\left(\boldsymbol{\mathrm{\zeta}}^{\text{new}}\right)$
The proposed learning-based DMPC is then outlined in Algorithm 2.
@@
14:     $\left(\boldsymbol{\mathrm{P}}_{i}^{k+1},\boldsymbol{\mathrm{Y}}_{i}^{k+1},\boldsymbol{\mathrm{w}}^{k+1}\right)=\textbf{{MABO}}\left(\boldsymbol{\mathrm{P}}_{i}^{k},\boldsymbol{\mathrm{Y}}_{i}^{k},\boldsymbol{\mathrm{w}}^{k}\right)$
@@
Algorithm 2 Coordinated-based Learning of DMPC
@@ IV-C Convergence Analysis of MABO
To investigate the convergence of the ADMM-based MABO, we rely on the standard ADMM convergence guarantees while accounting for the properties of the acquisition function $\alpha_{\text{EI}}$. Considering the optimization problem (35), the corresponding augmented Lagrangian is given by
$\sum_{i=1}^{m}\Big{(}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\boldsymbol{\mathrm{\lambda}}_{i}^{\top}(\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}})+\frac{\rho}{2}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}}\|^{2}\Big{)},$
The proof of convergence then proceeds in three steps below:
Step 1: Properties of $\alpha_{\text{EI}}$
The acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ is differentiable and satisfies Lipschitz continuity such that
@@
where $L$ is the Lipschitz constant. The boundedness and smoothness of $\alpha_{\text{EI}}$ then ensure that the local subproblem in (39) is solvable. Let us recall the acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ as:
$\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})=(J^{*}-\mu(\boldsymbol{\mathrm{\zeta}}_{i}))\Phi(Z)+\sigma(\boldsymbol{\mathrm{\zeta}}_{i})\phi(Z),$
@@
$Z=\frac{J^{*}-\mu(\boldsymbol{\mathrm{\zeta}}_{i})}{\sigma(\boldsymbol{\mathrm{\zeta}}_{i})}.$
Since $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ is a composition of differentiable functions, it is differentiable. The gradient of $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ with respect to $\boldsymbol{\mathrm{\zeta}}_{i}$ is obtained as
$\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})=\frac{\partial\alpha_{\text{EI}}}{\partial\mu}\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})+\frac{\partial\alpha_{\text{EI}}}{\partial\sigma}\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i}),$
where:
$\frac{\partial\alpha_{\text{EI}}}{\partial\mu}=-\Phi(Z),$
$\frac{\partial\alpha_{\text{EI}}}{\partial\sigma}=\phi(Z),$
and $\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i})$ are the gradients of the GP posterior mean and standard deviation, respectively. Both $\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i})$ are then bounded due to the smoothness of GP posterior predictions. Therefore, $\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ exists and is well-defined. To establish Lipschitz continuity of $\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$, we then compute the difference between gradients for two points $\boldsymbol{\mathrm{\zeta}}_{i}$ and $\boldsymbol{\mathrm{\zeta}}_{j}$ as
$\|\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})-\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{j})\|\leq\|\Delta_{\mu}\|+\|\Delta_{\sigma}\|,$
where:
@@
Using the smoothness of $\Phi(Z)$ and $\nabla\mu(\boldsymbol{\mathrm{\zeta}})$, there exists a constant $L_{\mu}$ such that
$\|\Delta_{\mu}\|\leq L_{\mu}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|.$
@@
Similarly, using the smoothness of $\phi(Z)$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}})$, there exists a constant $L_{\sigma}$ such that:
$\|\Delta_{\sigma}\|\leq L_{\sigma}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|.$
Combining the bounds for $\Delta_{\mu}$ and $\Delta_{\sigma}$, we then have that
$\|\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})-\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{j})\|\leq(L_{\mu}+L_{\sigma})\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|.$
Considering $L=L_{\mu}+L_{\sigma}$, the following then holds
$\|\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})-\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{j})\|\leq L\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|.$
@@
The acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ is then differentiable, and its gradient satisfies the Lipschitz continuity condition with constant $L=L_{\mu}+L_{\sigma}$.
Step 2: Boundedness of Lagrange Multipliers
Let us define the following residuals
$r_{1}^{k+1}=\|\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}\|,\quad r_{2}^{k+1}=\rho\|\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k}\|.$ (43)
The updates ensure that both the primal residual $r_{1}^{k+1}$ and dual residual $r_{2}^{k+1}$ converge to zero:
$r_{1}^{k+1},r_{2}^{k+1}\to 0\quad\text{as }k\to\infty.$ (44)
Let us recall the update rule of the dual variables $\boldsymbol{\mathrm{\lambda}}_{i}^{k}$ as
$\boldsymbol{\mathrm{\lambda}}_{i}^{k+1}=\boldsymbol{\mathrm{\lambda}}_{i}^{k}+\rho\left(\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}\right).$ (45)
The primal residual $r_{1}^{k+1}$ then measures how well the consensus constraint $\boldsymbol{\mathrm{\zeta}}_{i}=\boldsymbol{\mathrm{\bar{\zeta}}}$ is satisfied at iteration $k+1$. As $r_{1}^{k+1}\to 0$ (i.e., $\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}\to\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}$), the dual variable update then becomes stationary:
$\boldsymbol{\mathrm{\lambda}}_{i}^{k+1}=\boldsymbol{\mathrm{\lambda}}_{i}^{k}.$
The dual residual $r_{2}^{k+1}$ measures the change in the coordinating variable $\boldsymbol{\mathrm{\bar{\zeta}}}$ between successive iterations. As $r_{2}^{k+1}\to 0$ (i.e., $\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}\to\boldsymbol{\mathrm{\bar{\zeta}}}^{k}$), the updates for $\boldsymbol{\mathrm{\bar{\zeta}}}$ stabilize, indicating convergence. The ADMM updates then ensure that both residuals converge to zero:
$r_{1}^{k+1},\;r_{2}^{k+1}\to 0\quad\text{as }k\to\infty.$ (46)
We then have the following observations:
  • The convergence of $r_{1}^{k+1}\to 0$ implies that the consensus constraint $\boldsymbol{\mathrm{\zeta}}_{i}=\boldsymbol{\mathrm{\bar{\zeta}}}$ is satisfied.
  • The convergence of $r_{2}^{k+1}\to 0$ ensures that the coordinating variable $\boldsymbol{\mathrm{\bar{\zeta}}}$ stops changing between iterations.
  • Together, these residuals indicate that both the primal and dual updates stabilize, leading to the overall convergence of the ADMM scheme.
Although $\boldsymbol{\mathrm{\lambda}}_{i}^{k}$ are not explicitly included in the definitions of $r_{1}^{k+1}$ and $r_{2}^{k+1}$, they are inherently tied to the primal residuals because their updates depend on the difference $(\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1})$. Thus, the convergence of $r_{1}^{k+1}$ and $r_{2}^{k+1}$ ensures the stabilization of the dual variables.
Step 3: Decreasing Augmented Lagrangian
@@
The augmented Lagrangian $\mathcal{L}$ decreases monotonically under the updates:
@@
with equality only at the optimal solution.
Step 4: Convergence to Stationarity
Taking the gradient of the augmented Lagrangian with respect to $\boldsymbol{\mathrm{\zeta}}_{i}$ gives the necessary condition for the optimality of $\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}$:
@@
At convergence, we assume that $\boldsymbol{\mathrm{\zeta}}_{i}^{k}\to\boldsymbol{\mathrm{\zeta}}_{i}^{\star}$ and $\boldsymbol{\mathrm{\bar{\zeta}}}^{k}\to\boldsymbol{\mathrm{\bar{\zeta}}}^{\star}$, so the stationarity condition for $\boldsymbol{\mathrm{\zeta}}_{i}^{\star}$ becomes:

### v2

@@ IV-B Multi-Agent Bayesian Optimization
In this paper, we describe our approach to learn local MPC schemes in a distributed framework using BO. To this end, we leverage the BO in a multi-agent framework by introducing decomposable Multi-agent Bayesian Optimization (MABO). It should be noted that the use of non-myopic acquisition functions, as obtained in Proposition 1, can significantly increase the computational complexity. Moreover, the convergence analysis of ADMM-based MABO becomes more challenging when using non-myopic acquisition functions. In this paper, we then consider the myopic acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})$ to formulate a MABO algorithm. However, one may use the proposed non-myopic acquisition function for more complex objective functions in the proposed MABO since this look-ahead acquisition function $\tilde{\alpha}_{\text{EI}}(\boldsymbol{\mathrm{\theta}})$ can be simply replaced by $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})$ without changing the main structure of MABO.
@@
As observed in the formulation of the parametric local MPC scheme (33), each local value function depends on global parameters $\boldsymbol{\mathrm{\zeta}}:=\left\{\boldsymbol{\mathrm{\theta}}_{1},\ldots,\boldsymbol{\mathrm{\theta}}_{m}\right\}$ through optimal multipliers $\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\left(\boldsymbol{\mathrm{\zeta}}\right)$ (each agent has full access to all multipliers through a communication network between local agents). Consequently, each local parametric policy $\boldsymbol{\mathrm{\pi}}_{i}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)$ delivered by (33) is also affected by the global parameters $\boldsymbol{\mathrm{\zeta}}$ and its neighbor policies $\boldsymbol{\mathrm{\pi}}_{j}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)$ so that the corresponding local closed-loop control performance reads as $J_{i}\left(\boldsymbol{\mathrm{\zeta}}\right):=J_{i}\left(\boldsymbol{\mathrm{\pi}}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)\right)$ with
@@
An overview of the proposed learning-based DMPC scheme is illustrated in Fig. 1.
Figure 1: An overview of the proposed MABO-based DMPC.
@@
$J_{i}^{N}\left(\boldsymbol{\mathrm{\zeta}}\right)=\frac{1}{N}\sum_{k=0}^{N}L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}^{\boldsymbol{\mathrm{\zeta}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\hat{\boldsymbol{\mathrm{\mu}}}^{\star}\right)\right).$ (49)
We then implement a coordinated multi-agent learning approach by minimizing the global closed-loop performance as
@@
where $\rho>0$ is a constant penalty parameter and $\boldsymbol{\mathrm{\lambda}}=\left(\boldsymbol{\mathrm{\lambda}}_{1},\ldots,\boldsymbol{\mathrm{\lambda}}_{m}\right)$ denotes the set of dual variables (multipliers) associated with the equality constraint (51b). The local parameters are then obtained by the following local optimization problem
@@
where the coordinating variables $\boldsymbol{\mathrm{\Gamma}}_{i}^{k}:=(\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1},\boldsymbol{\mathrm{\lambda}}_{i}^{k})$ are updated by the following iterations
@@
To solve the problem (53), we then minimize the local acquisition function associated with $J_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i}\right)$ as
$\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}=\operatorname*{arg\,min}_{\boldsymbol{\mathrm{\zeta}}_{i}\in\mathcal{P}_{i}}\Big{(}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)\Big{)},$ (55)
where the penalty term $\Xi_{i}$ is
$\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)=(\boldsymbol{\mathrm{\lambda}}_{i}^{k})^{\top}\boldsymbol{\mathrm{\zeta}}_{i}+\frac{\rho}{2}\left\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}\right\|^{2}.$ (56)
Summary of the proposed multi-agent BO is given in Algorithm 1. The proposed learning-based DMPC is then outlined in Algorithm 2.
@@ 1  Input:
Coordinating variables $\boldsymbol{\mathrm{\Gamma}}_{i}^{k}:=(\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1},\boldsymbol{\mathrm{\lambda}}_{i}^{k})$
@@ 4  Output:
$\left(\boldsymbol{\mathrm{P}}_{i}^{k+1},\boldsymbol{\mathrm{Y}}_{i}^{k+1},\boldsymbol{\mathrm{\Gamma}}_{i}^{k+1}\right)=\textbf{{MABO}}\left(\boldsymbol{\mathrm{P}}_{i}^{k},\boldsymbol{\mathrm{Y}}_{i}^{k},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)$
@@
  3. Compute $\boldsymbol{\mathrm{\bar{\zeta}}}^{k+1}$ and $\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)$
  4. Find $\boldsymbol{\mathrm{\zeta}}_{i}^{\text{new}}=\operatorname*{arg\,min}_{\boldsymbol{\mathrm{\zeta}}_{i}\in\mathcal{P}_{i}}\Big{(}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\Xi_{i}\left(\boldsymbol{\mathrm{\zeta}}_{i},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)\Big{)}$
@@ 5. Evaluate $y_{i}^{\text{new}}=J_{i}^{N}\left(\boldsymbol{\mathrm{\zeta}}^{\text{new}}\right)$
14:     $\left(\boldsymbol{\mathrm{P}}_{i}^{k+1},\boldsymbol{\mathrm{Y}}_{i}^{k+1},\boldsymbol{\mathrm{\Gamma}}_{i}^{k+1}\right)=\textbf{{MABO}}\left(\boldsymbol{\mathrm{P}}_{i}^{k},\boldsymbol{\mathrm{Y}}_{i}^{k},\boldsymbol{\mathrm{\Gamma}}_{i}^{k}\right)$
@@
Algorithm 2 Coordinated learning of the DMPC schemes
@@ IV-C Convergence Analysis of MABO
To investigate the convergence of the ADMM-based MABO, we rely on the standard ADMM convergence guarantees while accounting for the properties of the acquisition function $\alpha_{\text{EI}}$.
Lemma 1.
The acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ defined in (39) is differentiable and satisfies Lipschitz continuity such that
@@
where $L$ is the Lipschitz constant.
Proof.
Consider the acquisition function $\alpha_{\text{EI}}$ as defined in (39). Since $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ is a composition of differentiable functions, it is differentiable. The gradient of $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ with respect to $\boldsymbol{\mathrm{\zeta}}_{i}$ is obtained as
$\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})=\frac{\partial\alpha_{\text{EI}}}{\partial\mu}\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})+\frac{\partial\alpha_{\text{EI}}}{\partial\sigma}\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i}),$ (58)
where $\frac{\partial\alpha_{\text{EI}}}{\partial\mu}=-\Phi(Z)$ and $\frac{\partial\alpha_{\text{EI}}}{\partial\sigma}=\phi(Z)$. $\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i})$ are the gradients of the GP posterior mean and standard deviation, respectively. Both $\nabla\mu(\boldsymbol{\mathrm{\zeta}}_{i})$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}}_{i})$ are then bounded due to the smoothness of GP posterior predictions. Therefore, $\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ exists and is well-defined. To establish the Lipschitz continuity of $\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$, we then compute the difference between gradients for two points $\boldsymbol{\mathrm{\zeta}}_{i}$ and $\boldsymbol{\mathrm{\zeta}}_{j}$ as
$\|\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})-\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{j})\|\leq\|\Delta_{\mu}\|+\|\Delta_{\sigma}\|,$ (59)
@@
Using the smoothness of $\Phi(Z)$ and $\nabla\mu(\boldsymbol{\mathrm{\zeta}})$, there exists a constant $L_{\mu}$ such that $\|\Delta_{\mu}\|\leq L_{\mu}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|$. Similarly, using the smoothness of $\phi(Z)$ and $\nabla\sigma(\boldsymbol{\mathrm{\zeta}})$, there exists a constant $L_{\sigma}$ such that $\|\Delta_{\sigma}\|\leq L_{\sigma}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|$. Combining the bounds for $\Delta_{\mu}$ and $\Delta_{\sigma}$, we then have that
$\|\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})-\nabla\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{j})\|\leq(L_{\mu}+L_{\sigma})\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\zeta}}_{j}\|.$ (60)
@@
Considering $L=L_{\mu}+L_{\sigma}$, the proof is complete. The acquisition function $\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})$ is then differentiable, and its gradient satisfies the Lipschitz continuity condition with the Lipschitz constant $L=L_{\mu}+L_{\sigma}$. ∎
Lemma 2.
The acquisition function $\alpha_{\text{EI}}$ defined in (39) is monotonically decreasing with respect to the exploration term $\alpha_{1}=\sigma\left(\boldsymbol{\mathrm{\theta}}\right)\in[0,1]$ and the exploitation term $\alpha_{2}=J^{\star}-\mu\left(\boldsymbol{\mathrm{\theta}}\right)$.
Proof.
Let us rewrite (39) as
$\alpha_{\text{EI}}(\cdot)=-\beta(\cdot),$ (61)
where $\beta(\cdot)=\alpha_{2}\Phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)+\alpha_{1}\phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)$. To show that $\alpha_{\text{EI}}(\cdot)$ is monotonically decreasing, one needs to show that $\beta(\cdot)$ is monotonically increasing. To this end, we take the derivative of $\beta(\cdot)$ w.r.t. $\alpha_{1}$ and $\alpha_{2}$ as
$\frac{\partial\beta}{\partial\alpha_{1}}=$ (62)
$-\alpha_{2}\Phi^{^{\prime}}\left(\frac{\alpha_{2}}{\alpha_{1}}\right)\left(\frac{\alpha_{2}}{\alpha_{1}^{2}}\right)+\phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)-\alpha_{1}\phi^{^{\prime}}\left(\frac{\alpha_{2}}{\alpha_{1}}\right)\left(\frac{\alpha_{2}}{\alpha_{1}^{2}}\right),$
$\frac{\partial\beta}{\partial\alpha_{2}}=\Phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)+\left(\frac{\alpha_{2}}{\alpha_{1}}\right)\Phi^{^{\prime}}\left(\frac{\alpha_{2}}{\alpha_{1}}\right)+\phi^{^{\prime}}\left(\frac{\alpha_{2}}{\alpha_{1}}\right).$ (63)
Let us consider the following properties of $\Phi$ and $\phi$:
$\phi^{^{\prime}}\left(x\right)=-x\phi\left(x\right),\quad\Phi^{^{\prime}}\left(x\right)=\phi\left(x\right).$ (64)
By leveraging the properties in (64), we can simplify the derivatives (62) and (63) such that the following holds:
$\frac{\partial\beta}{\partial\alpha_{1}}=\phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)>0,\quad\frac{\partial\beta}{\partial\alpha_{2}}=\Phi\left(\frac{\alpha_{2}}{\alpha_{1}}\right)>0.$ (65)
It then follows that $\alpha_{\text{EI}}(\cdot)=-\beta(\cdot)$ is monotonically decreasing since $\beta(\cdot)$ is monotonically increasing. ∎
Considering the optimization problem (51), the corresponding augmented Lagrangian is given by
$\sum_{i=1}^{m}\Big{(}\alpha_{\text{EI}}(\boldsymbol{\mathrm{\zeta}}_{i})+\boldsymbol{\mathrm{\lambda}}_{i}^{\top}(\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}})+\frac{\rho}{2}\|\boldsymbol{\mathrm{\zeta}}_{i}-\boldsymbol{\mathrm{\bar{\zeta}}}\|^{2}\Big{)}.$
@@
By Lemma 2, the augmented Lagrangian $\mathcal{L}$ decreases monotonically under the updates:
@@
with equality only at the optimal solution. By Lemma 1, $\nabla\alpha_{\text{EI}}$ then exists so that taking the gradient of the augmented Lagrangian with respect to $\boldsymbol{\mathrm{\zeta}}_{i}$ gives the necessary condition for the optimality of $\boldsymbol{\mathrm{\zeta}}_{i}^{k+1}$ as
@@
At convergence, we assume that $\boldsymbol{\mathrm{\zeta}}_{i}^{k}\to\boldsymbol{\mathrm{\zeta}}_{i}^{\star}$ and $\boldsymbol{\mathrm{\bar{\zeta}}}^{k}\to\boldsymbol{\mathrm{\bar{\zeta}}}^{\star}$, so the stationarity condition for $\boldsymbol{\mathrm{\zeta}}_{i}^{\star}$ becomes